OpenStudy (anonymous):

lim_{x rightarrow 1-} (sqrt{2x}\left| 1-x \right|)div(x ^{2}-1)

OpenStudy (saifoo.khan):

\[\lim_{x \rightarrow 1-} (\sqrt{2x}\left| 1-x \right|)\div(x ^{2}-1) \]

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

For 0<x<1 we have \[\frac{\sqrt{2x}|1-x|}{x^2-1}=-\frac{\sqrt{2x}(1-x)}{(x-1)(x+1)} = - \frac{\sqrt{2x}}{x+1} \] Now you can insert x=1, so the limit is \[-\sqrt{2}/2\]

OpenStudy (anonymous):

Oh sorry, erase the sign after the first equality! But that doesn't change anything.